The area of a triangular flag is 132 square centimeters. Its altitude is 2 centimeters longer than twice its base. Find the lengths of the altitude and the base. THANK YOU!!
Answer by ikleyn(52832)

.
You are given

0.5*A*B = 132    (1)   (area = 0.5 * (Altitude A) multiplied by (Base B) )
A = 2B +2.       (2)


===============>  by substitution (2) into (1)  ============>


(2B + 2)*B = 264,

2B^2 + 2B - 264 = 0  ====>

B^2 + B - 132 = 0

(B+12)*(B-11) = 0.


Only positive root makes sense:  B = 11 cm.


Altitude = 2B + 2 = 2*11 + 2 = 24 cm.